# Stimulating the Diffusion Model for Image Denoising via Adaptive Embedding and Ensembling

### v3

@@ 4.1 Overview
In this section, we propose a novel strategy to stimulate the diffusion model for image denoising (DMID), which mainly consists of an adaptive embedding method and an adaptive ensembling method. Our strategy performs a single inference by first embedding the input $y$ into an intermediate state $x_{N}$ of the diffusion model to match the following process:
@@
where $0\leq N<T$ is an intermediate timestep. We then generate a desirable denoised image by directly sampling from timestep $N$. The sampling strategy can be any existing approach, such as DDIM [51], DDRM [48], or DDNM [49] among others. In our experiments, we simply employ DDIM. The denoising process, achieved through direct sampling from timestep $N$, can be mathematically described as:
@@
where $x_{0}$ is the denoised output image. In addition, we employ our ensembling method to further reduce distortion. Specifically, we adjust the sampling times (referred to $S_{t}$) in one inference process and repeat the inference process several times (referred to $R_{t}$). The procedure for the inference process is described in Figure 3 and Algorithm 2. We will make explanations in later parts.
@@ 4.2 Adaptive Embedding
In this section, we will describe the procedure for embedding noisy images using eq. 7. The primary objective of our approach is to address general image denoising just by employing an unconditional diffusion model pre-trained for image synthesis. In essence, our embedding method is tasked with adaptively managing various types of noise and converting the noisy image into a state that the diffusion model can effectively deal with. To provide an overview, our embedding method follows a sequence of steps: it begins by transforming all types of noise into Gaussian noise, subsequently normalizes the image range to match the data range expected by diffusion models, and ultimately converts the image into the intermediate $x_{N}$ state of the diffusion model. In a broader sense, the function $Transform(y)$ in eq. 7 serves a dual purpose, functioning both as a noise transformation technique and a data normalizer. Initially, the function $Transform(y)$ transforms noisy images into a latent space where the Gaussian noise assumption is valid by existing noise transformation techniques [52, 53]. Subsequently, it adjusts the range of the latent noisy image to align with the data range of the diffusion model, which typically spans from -1 to 1. Following the $Transform(y)$ function, we multiply the latent noisy image by $\sqrt{\bar{\alpha}_{N}}$ to obtain the intermediate $x_{N}$ state of the diffusion model. The value of $N$ should be chosen carefully to achieve optimal denoising. After that, we conclude the embedding method. Furthermore, DMID performs denoising within this latent space and inverse transforms the latent denoised image, which is originally within the data range of the diffusion model, back to the original data range. In the subsequent parts, we will delve deeper into the specifics of the noise transformation technique and the significance of the parameter $N$. The noise transformation technique involves converting different types of noise into Gaussian noise. The primary goal of our approach is to address general image denoising, allowing it to handle various forms of image noise. Although we have successfully treated the diffusion model as a Gaussian denoiser, it is crucial to account for other noise types, such as real-world noise, which does not adhere to a Gaussian distribution. To surmount this challenge, we improve a noise transformation technique NN [53] to convert diverse noise types into Gaussian noise. In essence, noise transformation entails finding a latent image $z$ that exhibits correlation with the input noisy image $y$, while the noise in $z$ conforms to the assumption of additive white Gaussian noise (AWGN). NN [53] achieves this by training an encoder-decoder neural network using a single input noisy image for autoregressive modeling. The latent image $z$ is derived from the encoder and is expected to adhere to the AWGN assumption. We designate the encoder network of the VAE with parameters $\theta_{1}$ as $G_{\theta_{1}}$, the decoder network parameterized by $\theta_{2}$ as $F_{\theta_{2}}$, and the latent image as $z=G_{\theta_{1}}(y)\sim\mathcal{N}(x,\sigma^{2}I)$. The loss function of the noise transformation technique is presented as:
@@
where $x$ represents the clean image, $R(x)$ is a regularization function, and $\epsilon\sim\mathcal{N}(0,1)$. In the specific approach, three terms in the equation eq. 9 work in synergy. The first and second terms ensure data fidelity, while the final term serves as a regularization component. Together, these three terms prevent zero mapping and identity mapping. To facilitate unsupervised training using only the input noisy image $y$ without the clean image $x$, eq. 9 is optimized through ADMM [54]:
@@
where $\bar{x}$ is an estimation of the clean image $x$, $p$ is an auxiliary variable, $q$ is the dual variable and $\rho>0$ is a chosen constant. The subproblems are solved by alternating minimization since the clean image $x$ is not accessible during unsupervised training. Specifically eq. 11 is solved by updating the clean image estimation $\bar{x}$ and the network parameters $\theta_{1}$ and $\theta_{2}$ alternatively. In addition, eq. 12 is solved by its closed-form solution as:
@@
where $D$ is any existing Gaussian denoiser. We directly choose BM3D [22]. Instead of directly using the latent image $z=G_{\theta_{1}}(y)$, NN employs $\bar{x}+q/\rho$ as the final result after transformation. This result is a linear combination of the noisy image $y$ and the latent image $z$.
@@
[…]
Figure 4: Every 500 iterations, we employ SURE to determine whether to stop the noise transformation. A higher value of “-SURE” indicates better image quality. (a)-(d) illustrate the changes in PSNR and -SURE during the noise transformation process. In usual cases, the variations in -SURE align with those of PSNR.
The value of $N$ should be chosen carefully to achieve optimal denoising performance based on the noise level of the latent noisy image. Here, we multiply the latent noisy image by $\sqrt{\bar{\alpha}_{N}}$ to convert the image to the intermediate state $x_{N}$ within the diffusion model, as depicted in eq. 7. Subsequently, we generate a denoised image by directly sampling from $x_{N}$, as outlined in eq. 8. Since the diffusion model can be viewed as a denoiser, its denoising ability at timestep $N$ is ${\sqrt{1-\bar{\alpha}_{N}}}/{\sqrt{\bar{\alpha}_{N}}}$. To attain optimal denoising outcomes, aligning the denoising ability of the diffusion model at timestep $N$ with the noise level of the intermediate image $x_{N}$ is pivotal. Consequently, each value of $N$ timestep corresponds to a specific noise level $\sigma$ of the latent noisy image. When encountering a different noise level $\sigma_{1}$, we can readily identify an $N$ that corresponds to the noise level $\sigma_{2}$ closest to $\sigma_{1}$. For tasks such as Gaussian denoising, the noise level of the latent noisy image is known. In cases of unknown noise levels, established noise estimation techniques [58, 59] can be applied to address the issue.
@@ 4.3 Adaptive Ensembling
In this section, we will describe the procedure for the ensembling method. The primary objective of the ensembling method is to reduce distortion. In essence, our embedding method is tasked with adaptively adjusting distortion and perception based on specific requirements. The primary factor influencing distortion and perception is the stochasticity brought by the Gaussian noise in the additional noise item. In practice, excessive stochasticity can lead to significant distortion in the final denoised image. In theory, moderate stochasticity can facilitate convergence to better results. Thus our ensembling strategy involves adjusting the sampling times to control stochasticity in one inference process and repeating the inference process multiple times to constrain stochasticity and converge to higher-quality images. Firstly we constrain the stochasticity in one inference. The level of stochasticity is primarily determined by the number of sampling times since the additional noise item appears in each sampling process. A higher number of sampling times allows for more refinement and introduces more stochasticity, resulting in a more detailed image with improved perceptual quality. However, this also leads to a smaller weighted noisy image $y$ at each iteration, which reduces its controllability at the same time. Therefore, larger sampling times generally yield better perceptual quality but also result in larger distortion. To adjust distortion and perception, we recommend adjusting the number of sampling times according to the desired outcome. To reduce distortion, setting sampling times to be less than 10 is typically sufficient.
@@
Furthermore, we enhance the denoised images through the stochasticity in multiple inferences. The Gaussian noise in the additional noise item introduces stochasticity into the sampling process. This enables generating multiple clean images for the same noisy image when the sampling times are greater than one. The additional random Gaussian noise can, in terms of mathematical expectation, assist in converging to better results [43]. In addition, according to the Law of Large Numbers, as the sample size increases, the sample mean converges to the population expectation with probability 1. Here, the individual denoised images are regarded as samples and the average of multiple denoised images can approximate an overall expectation. The averaged outcome is enhanced and typically outperforms individual denoised images. Thus we repeat the inference process multiple times to generate multiple denoised images for the same noisy input, and the average of all candidate images is taken. This approach tends to reduce distortion and improve the final denoising results.
@@
In summary, for different noise levels, we first calculate the $N$ value in advance and perform inference. During the inference process, noisy images are first embedded to an intermediate state of the diffusion model and sampling is initiated. The sampling process can be adjusted based on specific requirements. If we aim to minimize distortion, we usually repeat the inference process several times to generate multiple candidate restored images, and finally take the average of all the images. Unlike previous methods that often insist on extremely high sampling times, our DMID strategy recommends maintaining the total sampling times consistent but reducing the per-inference sampling times by distributing them across multiple inferences. It is imperative to highlight that our contribution is pioneering in multiple facets. Firstly, We reframe the diffusion model as a Gaussian denoiser. This allows us to denoise in a single step or potentially optimize through multiple iterations. Secondly, We elucidate the embedding method and the noise-correlated sampling starting point for image denoising. This not only reduces the required sampling times in one inference but also enhances the quality of the results. Finally, we present the elucidation of how sampling times $S_{t}$ and repetition times $R_{t}$ impact distortion-based and perception-based quality. This helps reduce distortion and lays the foundation for further research in this area.
@@ 5.1 Implementation Details
We employ a pre-trained model from [18] which is trained on 256$\times$256 images from ImageNet [69], with full timesteps $T=1000$. Following [18], $\beta_{i}$ is constant increasing linearly from $0.0001$ to $0.02$, $\sigma_{t}$ is set to be $\gamma\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}$, and $\gamma=0.85$. For various levels of noise, we first pre-calculate the value of $N$ and conduct inference. Since we can traverse through the perception-distortion curve [19, 70], we present two variants of our method. The first variant, named “DMID-d”, achieves the least distortion and satisfactory perceptual quality. The second variant, named “DMID-p”, achieves the best perceptual quality and tolerable distortion. To achieve this, we need to determine the number of sampling times (referred to $S_{t}$) in an inference process and how many times will we repeat and get an average (referred to $R_{t}$). To ensure fairness, we set the full sampling times to be 1000 following [15, 46], that is $S_{t}*R_{t}=1000$ for the first variant. For the second variant, we set $R_{t}=1$ and the value of $S_{t}$ varies from 2 to 200 for different datasets and various noise levels. As described in Section 4.3, increasing the number of sampling times $S_{t}$ and the repetition times $R_{t}$ will reduce distortion while conversely improving perceptual quality. This is why we introduced the two variations. Further explanations can be found in Section 5.4.
@@ 5.2 Gaussian Image Denoising
$\sigma=50$ 28.21 / 0.8806 / 0.179 29.46 / 0.9081 / 0.145 29.61 / 0.9110 / 0.136 29.62 / 0.9105 / 0.132 29.92 / 0.9155 / 0.121 27.59 / 0.8722 / 0.087
$\sigma=100$ 25.34 / 0.8025 / 0.293 26.48 / 0.8482 / 0.252 26.61 / 0.8532 / 0.234 26.57 / 0.8501 / 0.231 27.01 / 0.8621 / 0.211 24.61 / 0.7987 / 0.156
$\sigma=150$ 23.69 / 0.7421 / 0.370 24.81 / 0.8032 / 0.331 24.93 / 0.8103 / 0.306 24.88 / 0.8043 / 0.307 25.39 / 0.8233 / 0.274 22.94 / 0.7465 / 0.210
$\sigma=200$ 22.52 / 0.6910 / 0.432 23.66 / 0.7673 / 0.393 23.78 / 0.7762 / 0.363 23.72 / 0.7674 / 0.368 24.28 / 0.7931 / 0.305 21.56 / 0.6932 / 0.259
ImageNet $\sigma=250$ 21.62 / 0.6448 / 0.478 22.80 / 0.7374 / 0.445 22.92 / 0.7479 / 0.409 22.84 / 0.7363 / 0.417 23.44 / 0.7666 / 0.360 20.87 / 0.6701 / 0.289
$\sigma=50$ 27.84 / 0.8844 / 0.226 28.51 / 0.8991 / 0.183 28.59 / 0.9011 / 0.177 28.63 / 0.9015 / 0.173 28.72 / 0.9031 / 0.169 26.63 / 0.8605 / 0.122
$\sigma=100$ 25.06 / 0.8090 / 0.365 25.76 / 0.8357 / 0.308 25.84 / 0.8389 / 0.291 25.86 / 0.8376 / 0.295 25.94 / 0.8399 / 0.279 23.94 / 0.7840 / 0.208
$\sigma=150$ 23.58 / 0.7555 / 0.449 24.32 / 0.7932 / 0.395 24.41 / 0.7980 / 0.367 24.41 / 0.7945 / 0.380 24.53 / 0.7989 / 0.354 22.47 / 0.7314 / 0.264
$\sigma=200$ 22.56 / 0.7131 / 0.514 23.36 / 0.7616 / 0.464 23.47 / 0.7682 / 0.426 23.44 / 0.7623 / 0.447 23.57 / 0.7683 / 0.380 21.37 / 0.6842 / 0.312
CBSD68 $\sigma=250$ 21.80 / 0.6773 / 0.558 22.64 / 0.7368 / 0.519 22.77 / 0.7451 / 0.475 22.72 / 0.7364 / 0.500 22.87 / 0.7459 / 0.443 20.77 / 0.6610 / 0.352
$\sigma=50$ 28.84 / 0.8921 / 0.247 29.86 / 0.9132 / 0.188 30.00 / 0.9153 / 0.185 30.02 / 0.9152 / 0.181 30.14 / 0.9174 / 0.179 27.90 / 0.8770 / 0.131
$\sigma=100$ 26.00 / 0.8206 / 0.387 27.16 / 0.8609 / 0.297 27.30 / 0.8642 / 0.287 27.27 / 0.8616 / 0.289 27.47 / 0.8668 / 0.271 25.31 / 0.8107 / 0.211
$\sigma=150$ 24.41 / 0.7670 / 0.475 25.69 / 0.8233 / 0.381 25.85 / 0.8281 / 0.363 25.78 / 0.8219 / 0.372 26.04 / 0.8325 / 0.341 24.06 / 0.7707 / 0.275
$\sigma=200$ 23.33 / 0.7228 / 0.543 24.65 / 0.7930 / 0.450 24.84 / 0.7999 / 0.422 24.74 / 0.7901 / 0.439 25.06 / 0.8043 / 0.368 22.99 / 0.7311 / 0.318
Kodak24 $\sigma=250$ 22.51 / 0.6830 / 0.589 23.89 / 0.7686 / 0.504 24.09 / 0.7772 / 0.469 23.94 / 0.7631 / 0.491 24.36 / 0.7821 / 0.427 22.44 / 0.7133 / 0.356
$\sigma=50$ 28.35 / 0.9078 / 0.180 30.04 / 0.935 / 0.140 30.29 / 0.9378 / 0.134 30.31 / 0.9378 / 0.132 30.50 / 0.9394 / 0.129 28.34 / 0.9112 / 0.092
$\sigma=100$ 25.59 / 0.8499 / 0.283 27.05 / 0.8914 / 0.233 27.25 / 0.8962 / 0.220 27.22 / 0.8945 / 0.218 27.49 / 0.8978 / 0.205 25.37 / 0.8560 / 0.158
$\sigma=150$ 23.98 / 0.8041 / 0.360 25.36 / 0.8570 / 0.300 25.57 / 0.8645 / 0.279 25.51 / 0.8604 / 0.281 25.82 / 0.8659 / 0.266 23.72 / 0.8148 / 0.209
$\sigma=200$ 22.85 / 0.7646 / 0.423 24.20 / 0.8279 / 0.354 24.41 / 0.8384 / 0.327 24.33 / 0.8321 / 0.334 24.70 / 0.8417 / 0.290 22.51 / 0.7816 / 0.252
McMaster $\sigma=250$ 21.96 / 0.7279 / 0.469 23.31 / 0.8031 / 0.402 23.54 / 0.8161 / 0.367 23.43 / 0.8077 / 0.380 23.79 / 0.8173 / 0.344 21.73 / 0.7553 / 0.290
@@
${}_{+}$ MIRNet MPRNet DeamNet NAFNet Uformer Restormer

### v5

@@ 4.1 Overview
In this section, we propose a novel strategy to stimulate the diffusion model for image denoising (DMID), which mainly consists of an adaptive embedding method and an adaptive ensembling method. In summary, we first perform the embedding method to construct the initial sampling state. Subsequently, we generate multiple denoised images using the pre-trained diffusion model through multiple inferences. Finally, we employ the ensembling method to reduce distortion.
The embedding method first transforms the input noisy image $y$ into a latent space where the Gaussian noise assumption is valid. Subsequently, the embedding method normalizes the image range to match the data range expected by diffusion models. Ultimately, the embedding method converts the latent image into the intermediate $x_{N}$ state of the diffusion model. The embedding method can be summarised as:
@@
where $0\leq N<T$ is an intermediate timestep. The function $Transform(y)$ serves as both a noise transformation technique and a data normalizer. Following the $Transform(y)$ function, we multiply the latent noisy image by $\sqrt{\bar{\alpha}_{N}}$ to obtain the intermediate $x_{N}$ state of the diffusion model. The value of $N$ should be chosen carefully according to the noise level of the latent noisy image to achieve optimal denoising. After that, we conclude the embedding method.
Furthermore, we then generate a desirable denoised image by directly sampling from timestep $N$. The sampling strategy can be any existing approach, such as DDIM [54], DDRM [50], or DDNM [51] among others. In our experiments, we simply employ DDIM. The denoising process, achieved through direct sampling from timestep $N$, can be mathematically described as:
@@
where $x_{0}$ is the denoised output image. DMID performs denoising within the latent space and, thus needs to inverse transform the latent denoised image, which is originally within the data range of the diffusion model, back to the original data range.
In addition, we employ our ensembling method to further reduce distortion. Specifically, we adjust the sampling times (referred to $S_{t}$) in one inference process and repeat the inference process several times (referred to $R_{t}$). Each inference on the same image corresponds to a Monte Carlo sampling, yielding $R_{t}$ denoised images. To mitigate stochasticity and distortion, we conduct a Minimum Mean Square Error (MMSE) averaging of all denoised images, ultimately obtaining the final result.
The whole procedure for the inference of our method is described in Figure 3 and Algorithm 2. We will delve into the procedures of the embedding method and ensembling method in the following sections.
Figure 4: The procedure for the embedding method. The embedding method first transforms the noise into Gaussian noise and subsequently normalizes the latent noisy image. After that, we search all the timestep $t$ to find a timestep $N$ to guarantee $\frac{\sqrt{1-\bar{\alpha}_{N}}}{\sqrt{\bar{\alpha}_{N}}}$ closest to $\sigma$. Ultimately, we multiply the noisy image by $\sqrt{\bar{\alpha}_{N}}$ to convert the image to the intermediate state $x_{N}$.
@@ 4.2 Adaptive Embedding
In this section, we will describe the procedure for the embedding method. As discussed before, the diffusion model is reframed as a Gaussian denoiser. The intermediate states within the diffusion model are adept at handling Gaussian noise with distinct levels. Consequently, our embedding method is tasked with transforming the noise within a noisy image into Gaussian noise and converting the image into a suitable intermediate state based on the noise level. The procedure of the embedding method is shown in Figure 4.
The noise transformation technique involves converting different types of noise into Gaussian noise. The primary goal of our approach is to address general image denoising, allowing it to handle various forms of image noise. Although we have successfully treated the diffusion model as a Gaussian denoiser, it is crucial to account for other noise types, such as real-world noise, which does not adhere to a Gaussian distribution. To surmount this challenge, we improve a noise transformation technique NN [57] to convert diverse noise types into Gaussian noise. In essence, noise transformation entails finding a latent image $z$ that exhibits correlation with the input noisy image $y$, while the noise in $z$ conforms to the assumption of additive white Gaussian noise (AWGN). NN [57] achieves this by training an encoder-decoder neural network using a single input noisy image for autoregressive modeling. The latent image $z$ is derived from the encoder and is expected to adhere to the AWGN assumption. We designate the encoder network of the VAE with parameters $\theta_{1}$ as $G_{\theta_{1}}$, the decoder network parameterized by $\theta_{2}$ as $F_{\theta_{2}}$, and the latent image as $z=G_{\theta_{1}}(y)\sim\mathcal{N}(x,\sigma^{2}I)$. The loss function of the noise transformation technique is presented as:
@@
where $x$ represents the clean image, $R(x)$ is a regularization function, and $\epsilon\sim\mathcal{N}(0,1)$. The derivation of Eq. 9 is given in the supplementary material. In the specific approach, three terms in the Eq. 9 work in synergy. The first and second terms ensure data fidelity, while the final term serves as a regularization component. Together, these three terms prevent zero mapping and identity mapping. To facilitate unsupervised training using only the input noisy image $y$ without the clean image $x$, Eq. 9 is optimized through ADMM [58]:
@@
where $\bar{x}$ is an estimation of the clean image $x$, $p$ is an auxiliary variable, $q$ is the dual variable and $\rho>0$ is a chosen constant. The subproblems are solved by alternating minimization since the clean image $x$ is not accessible during unsupervised training. Specifically Eq. 11 is solved by updating the clean image estimation $\bar{x}$ and the network parameters $\theta_{1}$ and $\theta_{2}$ alternatively. In addition, Eq. 12 is solved through the plug-and-play idea:
@@
where $D$ is any existing Gaussian denoiser. We choose BM3D [27]. Instead of directly using the latent image $z=G_{\theta_{1}}(y)$, NN employs a linear combination of the noisy image $y$ and the latent image $z$ as the final result after transformation.
@@
where $T$ is the transpose operator and $\mu$ is a small positive value. Every 500 iterations, we calculate the SURE of the latent image $z$. Smaller SURE values indicate higher image quality, and we stop the iterations when SURE begins to increase. It merits mentioning that our improved transformation technique requires no pre-training on external datasets or the use of the clean image $x$. The noise transformation process solely relies on the single input noisy image $y$.
The value of $N$ should be chosen carefully to achieve optimal denoising performance based on the noise level of the latent noisy image. For the latent noisy image with noise level $\sigma$, we search all the timestep $t$ to find a timestep $N$ to guarantee $\frac{\sqrt{1-\bar{\alpha}_{N}}}{\sqrt{\bar{\alpha}_{N}}}$ closest to $\sigma$. After that, we multiply the noisy image by $\sqrt{\bar{\alpha}_{N}}$ to convert the image to the intermediate state $x_{N}$. The reason is that we should align the denoising ability of the diffusion model at timestep $N$ with the noise level $\sigma$ of the latent noisy image to attain optimal denoising outcomes. And we consider the diffusion model as a denoiser, whose denoising ability at timestep $t$ is ${\sqrt{1-\bar{\alpha}_{t}}}/{\sqrt{\bar{\alpha}_{t}}}$, as discussed in Section 3.3. In practice, we calculate the corresponding level $\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}$ for each timestep $t$ in advance. For a latent noisy image with noise level $\sigma$, we readily identify an $N$ that corresponds to the level $\frac{\sqrt{1-\bar{\alpha}_{N}}}{\sqrt{\bar{\alpha}_{N}}}$ closest to $\sigma$. For instance, we choose $N=115$ for noise level $\sigma=\frac{2*50}{255}\approx 0.3922$, as ${\sqrt{1-\bar{\alpha}_{t}}}/{\sqrt{\bar{\alpha}_{t}}}$ closest to $\sigma$ at $t=115$ among all timestep $t$, as shown in Figure 4. For tasks such as Gaussian denoising, the noise level of the latent noisy image is known. In cases of unknown noise levels, established noise estimation techniques [62, 63] can be applied to address the issue.
@@ 4.3 Adaptive Ensembling
In this section, we will describe the procedure for the ensembling method. Our embedding method is tasked with adjusting distortion and perception based on specific requirements. The primary factor influencing distortion and perception is the stochasticity brought by the Gaussian noise in the additional noise item. In practice, excessive stochasticity can lead to significant distortion in the final denoised image. In theory, moderate stochasticity can facilitate convergence to better results [48]. Thus our ensembling strategy involves adjusting the sampling times to control stochasticity in one inference process and averaging the images obtained from multiple inferences to converge to higher quality.
Firstly we constrain the stochasticity in one inference. The level of stochasticity is primarily determined by the number of sampling times since the additional noise item appears in each sampling process. A higher number of sampling times $S_{t}$ allows for more refinement and introduces more stochasticity, resulting in a more detailed image with improved perceptual quality. However, this also leads to a smaller weighted noisy image $y$ at each iteration, which reduces its controllability at the same time. Therefore, larger sampling times generally yield better perceptual quality but also result in larger distortion. To adjust distortion and perception, we recommend adjusting the number of sampling times according to the desired outcome. To reduce distortion, setting sampling times to be less than 10 is typically sufficient.
@@
Furthermore, we enhance the denoised images through the stochasticity in multiple inferences. The Gaussian noise in the additional noise item introduces stochasticity into the sampling process. This enables generating multiple clean images for the same noisy image when the sampling times are greater than one. Thus, we propose an ensembling approach based on the Monte Carlo method to utilize the stochasticity of multiple inferences. Specifically, we repeat the inference process multiple times to generate multiple candidate denoised images for the same noisy input. We refer to the repetition times of inference as $R_{t}$, and the candidate denoised image as $c_{i}$. Additionally, we incorporate the Minimum Mean Square Error (MMSE) averaging [73] to further enhance the ensembling method. The final denoised image is expressed as
$x^{\text{MMSE}}=\frac{\sum_{i=1}^{R_{t}}c_{i}p(y|c_{i})}{\sum_{i=1}^{R_{t}}p(y|c_{i})}\ ,$ (17)
where $p(y|c_{i})\sim\mathcal{N}(c_{i},\sigma^{2}\mathbf{I})$ is calculated by the Gaussian noise model, $\sigma$ is the noise level of the noisy image, and $\mathcal{N}$ represents the Gaussian distribution. The derivation of Eq. 17 is given in the supplementary material. Inspired by our denoising prespective understanding, we reframe the diffusion model as a Gaussian denoiser and transform all noise to Gaussian noise. Thus we can leverage existing noisy images $y$ and the Gaussian noise model $p(y|x)$ to compute the observation likelihood $p(y|c_{i})$ for each candidate restored image $c_{i}$. This likelihood indicates how well the candidate restored image $c_{i}$ aligns with the denoised outcome of the noisy image $y$. Therefore, by weighting our predicted denoised results with the corresponding observation likelihood, we can further minimize stochasticity. According to the Law of Large Numbers, a higher number of repetition times $R_{t}$ yield better distortion-based quality.
@@
It is imperative to highlight that our contribution is pioneering in multiple facets. Firstly, we reframe the diffusion model as a Gaussian denoiser. This allows us to denoise in a single step or potentially optimize through multiple iterations. Secondly, we elucidate the embedding method and the noise-correlated sampling starting point for image denoising. This not only reduces the required sampling times in one inference but also enhances the quality of the results. Finally, we present the elucidation of how sampling times $S_{t}$ and repetition times $R_{t}$ impact distortion-based and perception-based quality. This helps reduce distortion and lays the foundation for further research in image denoising area.
@@ 5.1 Implementation Details
We employ a pre-trained model from [23] which is trained on 256$\times$256 images from ImageNet [74], with full timesteps $T=1000$. Following [23], $\beta_{t}$ is defined increasing linearly from $0.0001$ to $0.02$ for $t$ from $1$ to $1000$, $\sigma_{t}$ is set to be $\gamma\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}$, and $\gamma=0.85$.
Since our method can traverse through the perception-distortion curve [24, 75], we present two variants of our method. The first variant, named “DMID-d”, achieves the least distortion and satisfactory perceptual quality. The second variant, named “DMID-p”, achieves the best perceptual quality and tolerable distortion. To achieve this, we need to determine the number of sampling times (referred to $S_{t}$) in an inference process and how many times will we repeat and get an average (referred to $R_{t}$). For the first variant, we set the full sampling times $S_{t}*R_{t}$ to be 1000 following [15, 20] to ensure fairness. For the second variant, we set $R_{t}=1$ and the value of $S_{t}$ varies from 2 to 200 for different datasets and various noise levels. As described in Section 4.3, decreasing the number of sampling times $S_{t}$ and increasing the number of repetition times $R_{t}$ will reduce distortion while conversely improving perceptual quality. This is why we introduced the two variations. Further explanations can be found in Section 5.4.
@@ 5.2 Gaussian Image Denoising
In this section, we conduct Gaussian image denoising experiments on synthetic benchmark datasets.
Image denoising is commonly evaluated using the distortion-based PSNR metric with noise levels $\sigma=15,25,50$. However, our method is robust to much higher noise levels, excelling in both distortion-based and perception-based metrics. Hence, we perform classical Gaussian denoising experiments as well as robust Gaussian denoising experiments.
Subsequently, we will provide a detailed explanation of our comparative methods. DnCNN [7] and DRUNet [6] are considered classical methods, while Restormer [5] and ART [76] represent the current state-of-the-art approaches. For DnCNN [7] and DRUNet [6], we retrain them following their specified training details. However, it is worth noting that retraining Restormer [5] and ART [76] is a resource-intensive process. Their publicly available pre-trained models are not designed to handle noise levels with standard deviations exceeding 50. Therefore, following the approach of [6], we multiply by a constant to ensure that the standard deviation is 50.
@@
$\sigma=50$ 28.21 / 0.8806 / 0.179 29.46 / 0.9081 / 0.145 29.61 / 0.9110 / 0.136 29.62 / 0.9105 / 0.132 29.90 / 0.9157 / 0.114 27.59 / 0.8722 / 0.087
$\sigma=100$ 25.34 / 0.8025 / 0.293 26.48 / 0.8482 / 0.252 26.61 / 0.8532 / 0.234 26.57 / 0.8501 / 0.231 27.00 / 0.8626 / 0.201 24.61 / 0.7987 / 0.156
$\sigma=150$ 23.69 / 0.7421 / 0.370 24.81 / 0.8032 / 0.331 24.93 / 0.8103 / 0.306 24.88 / 0.8043 / 0.307 25.39 / 0.8236 / 0.263 22.94 / 0.7465 / 0.210
$\sigma=200$ 22.52 / 0.6910 / 0.432 23.66 / 0.7673 / 0.393 23.78 / 0.7762 / 0.363 23.72 / 0.7674 / 0.368 24.25 / 0.7915 / 0.295 21.56 / 0.6932 / 0.259
ImageNet $\sigma=250$ 21.62 / 0.6448 / 0.478 22.80 / 0.7374 / 0.445 22.92 / 0.7479 / 0.409 22.84 / 0.7363 / 0.417 23.44 / 0.7675 / 0.346 20.87 / 0.6701 / 0.289
$\sigma=50$ 27.84 / 0.8844 / 0.226 28.51 / 0.8991 / 0.183 28.59 / 0.9011 / 0.177 28.63 / 0.9015 / 0.173 28.69 / 0.9029 / 0.162 26.63 / 0.8605 / 0.122
$\sigma=100$ 25.06 / 0.8090 / 0.365 25.76 / 0.8357 / 0.308 25.84 / 0.8389 / 0.291 25.86 / 0.8376 / 0.295 25.96 / 0.8413 / 0.283 23.94 / 0.7840 / 0.208
$\sigma=150$ 23.58 / 0.7555 / 0.449 24.32 / 0.7932 / 0.395 24.41 / 0.7980 / 0.367 24.41 / 0.7945 / 0.380 24.54 / 0.8001 / 0.361 22.47 / 0.7314 / 0.264
$\sigma=200$ 22.56 / 0.7131 / 0.514 23.36 / 0.7616 / 0.464 23.47 / 0.7682 / 0.426 23.44 / 0.7623 / 0.447 23.57 / 0.7686 / 0.392 21.37 / 0.6842 / 0.312
CBSD68 $\sigma=250$ 21.80 / 0.6773 / 0.558 22.64 / 0.7368 / 0.519 22.77 / 0.7451 / 0.475 22.72 / 0.7364 / 0.500 22.88 / 0.7462 / 0.455 20.77 / 0.6610 / 0.352
$\sigma=50$ 28.84 / 0.8921 / 0.247 29.86 / 0.9132 / 0.188 30.00 / 0.9153 / 0.185 30.02 / 0.9152 / 0.181 30.13 / 0.9174 / 0.172 27.90 / 0.8770 / 0.131
$\sigma=100$ 26.00 / 0.8206 / 0.387 27.16 / 0.8609 / 0.297 27.30 / 0.8642 / 0.287 27.27 / 0.8616 / 0.289 27.50 / 0.8682 / 0.274 25.31 / 0.8107 / 0.211
$\sigma=150$ 24.41 / 0.7670 / 0.475 25.69 / 0.8233 / 0.381 25.85 / 0.8281 / 0.363 25.78 / 0.8219 / 0.372 26.08 / 0.8336 / 0.348 24.06 / 0.7707 / 0.275
$\sigma=200$ 23.33 / 0.7228 / 0.543 24.65 / 0.7930 / 0.450 24.84 / 0.7999 / 0.422 24.74 / 0.7901 / 0.439 25.08 / 0.8054 / 0.382 22.99 / 0.7311 / 0.318
Kodak24 $\sigma=250$ 22.51 / 0.6830 / 0.589 23.89 / 0.7686 / 0.504 24.09 / 0.7772 / 0.469 23.94 / 0.7631 / 0.491 24.40 / 0.7853 / 0.446 22.44 / 0.7133 / 0.356
$\sigma=50$ 28.35 / 0.9078 / 0.180 30.04 / 0.9350 / 0.140 30.29 / 0.9378 / 0.134 30.31 / 0.9378 / 0.132 30.51 / 0.9396 / 0.124 28.34 / 0.9112 / 0.092
$\sigma=100$ 25.59 / 0.8499 / 0.283 27.05 / 0.8914 / 0.233 27.25 / 0.8962 / 0.220 27.22 / 0.8945 / 0.218 27.57 / 0.8990 / 0.206 25.37 / 0.8560 / 0.158
$\sigma=150$ 23.98 / 0.8041 / 0.360 25.36 / 0.8570 / 0.300 25.57 / 0.8645 / 0.279 25.51 / 0.8604 / 0.281 25.89 / 0.8672 / 0.267 23.72 / 0.8148 / 0.209
$\sigma=200$ 22.85 / 0.7646 / 0.423 24.20 / 0.8279 / 0.354 24.41 / 0.8384 / 0.327 24.33 / 0.8321 / 0.334 24.68 / 0.8392 / 0.303 22.51 / 0.7816 / 0.252
McMaster $\sigma=250$ 21.96 / 0.7279 / 0.469 23.31 / 0.8031 / 0.402 23.54 / 0.8161 / 0.367 23.43 / 0.8077 / 0.380 23.81 / 0.8174 / 0.359 21.73 / 0.7553 / 0.290
@@
DANet+ MIRNet MPRNet DeamNet NAFNet Uformer Restormer